# Optimal Task Order for Continual Learning of Multiple Tasks

### v1

@@ 5.2 Numerical results
We estimated the performance of the periphery-to-core rule and max-path rule in task-incremental continual learning using Fashion-MNIST [51], CIFAR-10, and CIFAR-100 dataset [52] (see Appendix C for the details). For the Fashion-MNIST and CIFAR-10 datasets, we randomly generated five binary image classification tasks by dividing 10 labels into 5 pairs without replacement. In the case of CIFAR-100, we selected 10 labels out of 100 labels randomly and generated 5 binary classifications. For Fashion-MNIST, we trained a multi-layered perceptron with two hidden layers, while for CIFAR-10/100, we used a convolutional neural network with two convolutional layers and one dense layer, to explore robustness against the model architecture.
We found that the final performance was modulated by the estimated average similarity among tasks, $\bar{\rho}=\tfrac{1}{P(P-1)}\sum_{\mu\neq\nu}\rho_{\mu\nu}$, we thus plotted the performance of each task-order rule as a function of the average similarity (Fig. 5). In all three settings, we found that the periphery-to-core rule robustly outperforms the core-to-periphery rule and average performance over random ordering (Fig. 5a,c,e). Similarly, the max-path rule outperformed both the min-path rule and the random ordering (Fig. 5b,d,e; see also Bell and Lawrence [20]). Moreover, we observed consistent results under a continual learning of a multi-class classification (Fig.7a and b in the Appendix). The periphery-to-core rule outperformed the max-path rule on average, bu the difference was small (red lines in Fig. 5 top vs bottom). When we increased the number of binary classification tasks from 3 to 7 using CIFAR-100, the performance advantage periphery-to-core over core-to-periphery increased (Fig. 5g) as expected from the linear model (Fig. 6c in the Appendix). This was not evident for max-path and min-path rules potentially because the difference was already high under $P=3$ (Fig. 5h). These results suggest the robustness of the task order optimization principles found in our simple analysis.
@@ 6 Discussion
In this work, we derived a simple analytical expression to explain how task similarity and ordering influence continual learning performance in a linear model with latent structure. Based on this result, we proposed two principles for task order optimization: the periphery-to-core rule and the max-path rule, the latter of which was predicted by Bell and Lawrence [20]. We validated these principles in task-incremental continual image classification tasks using both multi-layer perceptrons and convolutional neural networks. Thus, this work proposes basic principles for task order optimization in the context of continual learning for multiple tasks.
@@
Acknowledgment
This work was supported by Washington University in St Louis.
@@ A.1 Model setting
Below, we analyze task-order dependence of continual learning using linear teacher-student models with a latent factor. In teacher-student models, the generative model of the task parameterized explicitly by the teacher model, making the learning dynamics and the performance analytically tractable [31, 36, 32]. Here, the generative model for input $\bm{x}\in\mathbb{R}^{N_{x}}$ and the target output $\bm{y}\in\mathbb{R}^{N_{y}}$ is constructed as
@@
The second equality follows from the Gaussianity of $\bm{s}$. We consider task-incremental continual learning [50] where the network is trained for task $\mu=1,...,P$ in sequence. During the training for the $\mu$-th task, weight $W$ is updated by gradient descent on error $\epsilon_{\mu}$:
@@ A.2 The weight after continual learning of $P$ tasks
To further investigate how task similarity impacts continual learning performance, below we focus on the large $N_{x}$ regime, and analyze the learning behavior at $\tfrac{N_{s}}{N_{x}}\to 0$ limit. This assumption of the presence of low-dimensional latent factor is consistent with many real-world datasets [48, 49]. If $A_{\mu}$ is a very-tall random matrix (i.e., $N_{x}\gg N_{s}$), pseudo-inverse $A_{\mu}^{+}$ is approximated by a scaled transpose $\gamma A_{\mu}^{T}$, and $U_{\mu}U_{\mu}^{T}$ approximately follows $U_{\mu}U_{\mu}^{T}\approx\gamma A_{\mu}A_{\mu}^{T}$, where $\gamma=\tfrac{N_{s}}{N_{x}}$ (see Appendix A.4). Thus, we have
@@
Note that, if $C^{in}=I$, the error is zero. This is consistent with previous results showing that in the absence of overlap between tasks, continual learning doesn’t suffer from forgetting [5, 6, 53]. Additionally, Eq. 33 requires $P\ll\tfrac{N_{x}}{N_{s}}$ (see Appendix A).4 below), thus the obtained expression doesn’t hold when the number of tasks is comparable to the network size.
@@ A.4 Expectation over random correlated matrices $\{A_{\mu}\}$
$\delta_{j_{0},j_{1},...,j_{k}}$ in the third line is the Kronecker delta function that returns 1 if $j_{0}=j_{1}=...=j_{k}$, otherwise returns 0. The third line follows from Isserlis’ theorem, which states that the expectation over multivariate normal variables can be decomposed into summation over all pair-wise partitions [54]. In the equation above, the partition that pairs neighboring matrices takes $\mathcal{O}(N_{x}^{k+1})$ value, while all other partitions yield $\mathcal{O}(N_{x}^{k})$ value at most because of indices mismatch. Note that, the number of second order terms depends on $P$, as suggested by the summation over $l$ in the third line. Thus, we expect that our theory hold only when $P$ satisfies $P\ll\tfrac{N_{x}}{N_{s}}$.
@@ C.3 Convolutional and multi-layered non-linear neural networks for image classification
The networks were trained by minimizing the cross-entropy loss using the Adam optimizer with a learning rate of $10^{-3}$ for five epochs per task. We set the batch size to 4 due to GPU memory constraints. The models were implemented using Flax [55], a JAX neural network library, and were trained on NVIDIA Tesla V100 GPUs
@@ C.4 Estimation of task similarity
In the main text, we inferred similarity between two tasks A and B using zero-shot transfer performance between the two tasks. Previous work on linear model indicates that, if the output similarity is one, the pairwise transfer performance $\Delta\epsilon_{TF}[\nu\to\mu]\equiv\epsilon_{\mu}[W_{\nu}]-\epsilon_{\mu}[W_{o}]$ is written as [46]
@@
Motivated by this relationship, we defined similarity between tasks in general nonlinear networks by Eq. 16. A similar approach was implemented in [17]. Notably, this method only requires inputs/outputs of the trained network, and thus applicable to situation where the model details are inaccessible (e.g., human and animal brains, closed-LLM).
We implemented the evaluation the zero-shot transfer performance from task A to B as follows: First, we trained a network on task A for 5 epochs from a random initialization using the Adam optimizer on the cross-entropy loss with learning rate $10^{-3}$ as above. We then measured the cross-entropy loss on the test dataset of task B, $\epsilon_{B}[W_{A}]$. To normalize the accuracy, we divided the obtained loss by the loss on task B under a label shuffling, $\epsilon_{B,sf}[W_{A}]$. The resultant value $\frac{\epsilon_{B}[W_{A}]}{\epsilon_{B,sf}[W_{A}]}$ characterizes how well the network transfer to task B compared to a random task with the same input statistics. Since evaluating similarity using the transfer performance on the test data may potentially introduce bias, in Fig. 7c and d, we estimated the transfer performance $\Delta\epsilon_{B}[W_{A}]$ using the training dataset for task $B$. Even in this setting, we found results nearly identical to those in Fig. 5c and d, confirming the robustness of our findings with respect to the details of the similarity evaluation method.

### v2

@@ 5.2 Numerical Results
We estimated the performance of the periphery-to-core rule and max-path rule in task-incremental continual learning using Fashion-MNIST (Xiao et al., 2017), CIFAR-10, and CIFAR-100 dataset (Krizhevsky et al., 2009) (see Appendix C for the details). For the Fashion-MNIST and CIFAR-10 datasets, we randomly generated five binary image classification tasks by dividing 10 labels into 5 pairs without replacement. In the case of CIFAR-100, we selected 10 labels out of 100 labels randomly and generated 5 binary classifications. For Fashion-MNIST, we trained a multi-layered perceptron with two hidden layers, while for CIFAR-10/100, we used a convolutional neural network with two convolutional layers and one dense layer, to explore robustness against the model architecture.
We found that the final performance was modulated by the estimated average similarity among tasks, $\bar{\rho}=\tfrac{1}{P(P-1)}\sum_{\mu\neq\nu}\rho_{\mu\nu}$, we thus plotted the performance of each task-order rule as a function of the average similarity (Fig. 5). In all three settings, we found that the periphery-to-core rule robustly outperforms the core-to-periphery rule and average performance over random ordering (Fig. 5a,c,e). Similarly, the max-path rule outperformed both the min-path rule and the random ordering (Fig. 5b,d,e; see also Bell & Lawrence (2022)). Moreover, we observed consistent results under a continual learning of a multi-class classification (Fig.7a and b). The periphery-to-core rule outperformed the max-path rule on average, but the difference was small (red lines in Fig. 5 top vs bottom). When we increased the number of binary classification tasks from 3 to 7 using CIFAR-100, the performance advantage periphery-to-core over core-to-periphery increased (Fig. 5g) as expected from the linear model (Fig. 6c). This was not evident for max-path and min-path rules potentially because the difference was already high under $P=3$ (Fig. 5h).
We also investigated the inference of task similarity and ordering from a small subset of training data. This extension is crucial, as it demonstrates the practical relevance of our theory to real-world machine learning settings where full access to all training data upfront is often unrealistic (in contrast, when complete data is available, naive multi-task learning may suffice). When we reduced the number of training samples used to estimate task similarity across tasks in CIFAR-10, the relative advantage of both the periphery-to-core rule and the max-path rule over a random task order remained robust. Even when only 1% of the training data was used for estimating task similarity, we observed a performance gain comparable to that in the full data scenario (Panels c and d vs. g and h in Fig. 9). However, when the amount of data was reduced to 0.1% (approximately 10 samples per task), the performance gain became non-significant. We observed similar trends with both the Fashion-MNIST and CIFAR-100 datasets, although the results for CIFAR-100 were less robust, particularly in the negative similarity regime (Fig. 10). These results suggest the robustness of the task order optimization principles found in our simple analysis.
@@ 6 Discussion
In this work, we derived a simple analytical expression to explain how task similarity and ordering influence continual learning performance in a linear model with latent structure. Based on this result, we proposed two principles for task order optimization: the periphery-to-core rule and the max-path rule, the latter of which was predicted by Bell & Lawrence (2022). We validated these principles in task-incremental continual image classification tasks using both multi-layer perceptrons and convolutional neural networks. Thus, this work proposes basic principles for task order optimization in the context of continual learning for multiple tasks.
@@
Acknowledgements
This work was partially supported by McDonnell Center for Systems Neuroscience.
Impact Statement
Task order optimization for continual learning of multiple tasks may potentially contribute beyond the field of machine learning from school curriculum design (Rafferty et al., 2016; Zhu et al., 2018) to animal training protocol in neuroscience experiments (Krueger & Dayan, 2009). Nevertheless, due to theoretical nature of this work, there are no specific societal consequence that we feel must be highlighted here.
@@ A.1 Model Setting
Below, we analyze task-order dependence of continual learning using linear teacher-student models with a latent factor. In teacher-student models, the generative model of the task parameterized explicitly by the teacher model, making the learning dynamics and the performance analytically tractable (Gardner & Derrida, 1989; Saad & Solla, 1995; Zdeborová & Krzakala, 2016). Here, the generative model for input $\bm{x}\in\mathbb{R}^{N_{x}}$ and the target output $\bm{y}\in\mathbb{R}^{N_{y}}$ is constructed as
@@
The second equality follows from the Gaussianity of $\bm{s}$. We consider task-incremental continual learning (Van de Ven & Tolias, 2019) where the network is trained for task $\mu=1,...,P$ in sequence. During the training for the $\mu$-th task, weight $W$ is updated by gradient descent on error $\epsilon_{\mu}$:
@@ A.2 The Weight after Continual Learning of $P$ Tasks
To further investigate how task similarity impacts continual learning performance, below we focus on the large $N_{x}$ regime, and analyze the learning behavior at $\tfrac{N_{s}}{N_{x}}\to 0$ limit. This assumption of the presence of low-dimensional latent factor is consistent with many real-world datasets (Yu et al., 2017; Cohen et al., 2020). If $A_{\mu}$ is a very-tall random matrix (i.e., $N_{x}\gg N_{s}$), pseudo-inverse $A_{\mu}^{+}$ is approximated by a scaled transpose $\gamma A_{\mu}^{T}$, and $U_{\mu}U_{\mu}^{T}$ approximately follows $U_{\mu}U_{\mu}^{T}\approx\gamma A_{\mu}A_{\mu}^{T}$, where $\gamma=\tfrac{N_{s}}{N_{x}}$ (see Appendix A.4). Thus, we have
@@
Note that, if $C^{in}=I$, the error is zero. This is consistent with previous results showing that in the absence of overlap between tasks, continual learning doesn’t suffer from forgetting (Ramasesh et al., ; Lee et al., 2021; Peng et al., 2023). Additionally, Eq. 33 requires $P\ll\tfrac{N_{x}}{N_{s}}$ (see Appendix A).4 below), thus the obtained expression doesn’t hold when the number of tasks is comparable to the network size.
@@ A.4 Expectation over Random Correlated Matrices $\{A_{\mu}\}$
$\delta_{j_{0},j_{1},...,j_{k}}$ in the third line is the Kronecker delta function that returns 1 if $j_{0}=j_{1}=...=j_{k}$, otherwise returns 0. The third line follows from Isserlis’ theorem, which states that the expectation over multivariate normal variables can be decomposed into summation over all pair-wise partitions (Helias & Dahmen, 2020). In the equation above, the partition that pairs neighboring matrices takes $\mathcal{O}(N_{x}^{k+1})$ value, while all other partitions yield $\mathcal{O}(N_{x}^{k})$ value at most because of indices mismatch. Note that, the number of second order terms depends on $P$, as suggested by the summation over $l$ in the third line. Thus, we expect that our theory hold only when $P$ satisfies $P\ll\tfrac{N_{x}}{N_{s}}$.
@@ C.2 Generation of Input Similarity Matrices Having Simple Graph Structures
Figure 9: Task order preference estimated from various fraction of training data under CIFAR-10 task. For instance, in $0.1\%$ data results depicted in panels a and b, we first used $0.1\%$ of training data ($\sim 10$ images) for training the network with one task and used $0.1\%$ of training data from another task to evaluate the zero-shot transfer performance. We then estimated the task similarity as before. Panels g and h are the same with panels c and d in Fig. 7. We used the same set of parameters with the results depicted in the main figure (see Appendix C for details).
@@ C.3 Convolutional and Multi-layered Non-linear Neural Networks for Image Classification
The networks were trained by minimizing the cross-entropy loss using the Adam optimizer with a learning rate of $10^{-3}$ for five epochs per task. We set the batch size to 4 due to GPU memory constraints. The models were implemented using Flax (Heek et al., 2024), a JAX neural network library, and were trained on NVIDIA Tesla V100 GPUs
@@
Figure 10: Task order preference in continuous image classification tasks where the task similarity was estimated from 1% of training data, as opposed to Fig. 5 where 100% of data was used. a–f) Continual learning performance, defined as the average test accuracy across all the tasks after learning, under various task orders. g, h) The ratio of task sets where the periphery-to-core rule outperforms the core-to-periphery rule (g), and where the max-path rule outperforms the min-path rule (h), under CIFAR-100 with different numbers of tasks ($P=3,5,7$). Panels c and d were replicated from Fig. 9 for completeness. Note that task-order effects were smaller in CIFAR-100 because the number of training images per label is fewer in CIFAR-100 than in CIFAR-10 and Fashion-MNIST.
@@ C.4 Estimation of Task Similarity
In the main text, we inferred similarity between two tasks A and B using zero-shot transfer performance between the two tasks. Previous work on linear model indicates that, if the output similarity is one, the pairwise transfer performance $\Delta\epsilon_{TF}[\nu\to\mu]\equiv\epsilon_{\mu}[W_{\nu}]-\epsilon_{\mu}[W_{o}]$ is written as (Hiratani, )
@@
Motivated by this relationship, we defined similarity between tasks in general nonlinear networks by Eq. 16. A similar approach was implemented in (Lad et al., 2009). Notably, this method only requires inputs/outputs of the trained network, and thus applicable to situation where the model details are inaccessible (e.g., human and animal brains, closed-LLM).
We implemented the evaluation the zero-shot transfer performance from task A to B as follows: First, we trained a network on task A for 5 epochs from a random initialization using the Adam optimizer on the cross-entropy loss with learning rate $10^{-3}$ as above. We then measured the cross-entropy loss on the test dataset of task B, $\epsilon_{B}[W_{A}]$, where $W_{A}$ represents the weight after 5 epochs of training on task A. To normalize the accuracy, we divided the obtained loss by the loss on task B under a label shuffling, $\epsilon_{B,sf}[W_{A}]$. The resultant value $\frac{\epsilon_{B}[W_{A}]}{\epsilon_{B,sf}[W_{A}]}$ characterizes how well the network transfer to task B compared to a random task with the same input statistics. Since evaluating similarity using the transfer performance on the test data may potentially introduce bias, in Fig. 7c and d, we estimated the transfer performance $\Delta\epsilon_{B}[W_{A}]$ using the training dataset for task $B$. Even in this setting, we found results nearly identical to those in Fig. 5c and d, confirming the robustness of our findings with respect to the details of the similarity evaluation method.
In the task similarity and order estimation from sparse data shown in Figs. 9 and 10, we estimated task similarity in the same manner as described above, but using subsampled training data. For example, in the 1% data scenario, we used 1% of the training data for task A to calculate the weights after learning task A, denoted as $W_{A}$. As before, we trained the network using the Adam optimizer on the cross-entropy loss for 5 epochs from random initialization except that we only used 1% of training data at each epoch. We then estimated the transfer performance to task B, $\epsilon_{B}[W_{A}]$, using 1% of the training data for task B. The task similarity between tasks A and B was then computed using Eq. 16, as in the full data setting.